On January 1, 2019, Romero Company purchased equipment for $265,000.
The equipment was assigned a $7,000 residual value and a 16-year
life. Romero Company will use the straight-line method to depreciate
the equipment. On January 1, 2028, Romero Company spent $29,725 to
overhaul the equipment. This capital expenditure resulted in Romero
Company changing the life of the equipment from 16 years to 22 years
and adjusting the residual value to be $4,000 at the end of the 22
years.

Calculate the depreciation expense recorded on the equipment for 2028.
Answer #1

Depreciation as per SLM= (Purchase price-residual value)/ Useful life

so depreciation for 2019- 2027 would be= (265000-7000)/16= $ 16,125

After 9yrs, the wdv would be= 265000-(16125*9)= $119,875

After 9 yrs, the company incurred additional capital cost $ 29,725. The life of asset would be 13 yrs more with residual value $ 4000.

Now the depreciation would be= [(119,875+29,725)-4000]/13= $11,200

The depreciation in 2028 would be $ 11,200.